Give a rough estimate (by rounding off to nearest hundreds) and also a closer estimate (by rounding off to nearest tens): 
108734 – 47599

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

108734–47599
Rounding off to nearest hundreds
108734–47599 = 108700–47600 = 61100
Rounding off to nearest tens
108734–47599 = 108730–47600 = 61130
​​
Thanks